# Operator World Models for Reinforcement Learning

### v1

@@ A.2 Markov operators and their properties
Recall that since they are Hilbert spaces ${\mathcal{F}}\cong{\mathcal{F}}^{*}$ and ${\mathcal{G}}\cong{\mathcal{G}}^{*}$ are isometric to their dual and therefore we can interpret any $f\in{\mathcal{F}}$ as the function $f(\cdot)=\left\langle{f},{\varphi(\cdot)}\right\rangle$ with some abuse of notation, where clear from context. By Asm. 1 we have that ${T}|_{\mathcal{F}}$ takes values in ${\mathcal{G}}$. This means that $({T}|_{\mathcal{F}}f)\in{\mathcal{G}}$ or, in other words
$\left\langle{{T}|_{\mathcal{F}}f},{\psi(x,a)}\right\rangle$ $=({T}|_{\mathcal{F}}f)(x,a)$
@@
$\left\langle{f},{({T}|_{\mathcal{F}})^{*}\psi(x,a)}\right\rangle=\left\langle{f},{\int\varphi(x^{\prime})\leavevmode\nobreak\ \tau(dx^{\prime}|x,a)}\right\rangle.$
@@
We note that the result can be extended to the setting where ${T}|_{{\mathcal{F}}}({\mathcal{F}})\subseteq{\mathcal{G}}$, namely the image of ${T}|_{{\mathcal{F}}}$ is contained in ${\mathcal{G}}$, namely a sort of $({\mathcal{F}},{\mathcal{G}})$-compatibility for the transition operator (see Definition 1).
@@ A.3 The operatorial formulation of RL
According to the operatorial characterization in 7, the action value function of a policy $\pi$ is directly related to the action of the associated policy operator ${P}$. To highlight this relation, we will adopt the following notation:
@@
    $q({P})=\left({Id}-\gamma{T}{P}\right)^{-1}r.$ (A.5)
@@
    $J({P})=\left\langle{{P}\left({Id}-\gamma{T}{P}\right)^{-1}r},{\nu}\right\rangle=\left\langle{{P}q({P})},{\nu}\right\rangle=\left\langle{v({P})},{\nu}\right\rangle.$ (A.7)
  • State visitation (or State occupancy) measure. By the characterization of the adjoints of ${P}$ and ${T}$ (see discussion in Sec. 3 we can represent the evolution of a state distribution $\nu_{t}$ at time $t$ to the next state distribution as $\nu_{t+1}={T}^{*}{P}^{*}\nu_{t}$. Applying this relation recursively, we recover the state visitation probability associated to the starting state distribution $\nu_{0}=\nu\in{\mathcal{M}}_{1}^{+}(\mathcal{X})$, the MDP with transition ${T}$ and the policy ${P}$ as
    $d_{\nu}({P})=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}({T}^{*}{P}^{*})^{t}\nu=(1-\gamma)\left({Id}-\gamma{P}{T}\right)^{-*}\nu,$ (A.8)
    where the $(1-\gamma)\gamma^{t}$ is a normalizing factor to guarantee that the series corresponds to a convex combination of the probability distributions $\nu_{t}$, hence guaranteeing $d_{\nu}({P})$ to be well-defined (namely it belongs to ${\mathcal{M}}_{1}^{+}(\mathcal{X})$).
@@
Algebraic manipulation of the cumulative expected reward $J({P})$ implies
$\begin{split}J({P})&=\left\langle{{P}\left({Id}-\gamma{T}{P}\right)^{-1}r},{\nu}\right\rangle=\left\langle{{P}r},{\left({Id}-\gamma{P}{T}\right)^{-*}\nu}\right\rangle=\frac{1}{1-\gamma}\left\langle{{P}r},{d_{\nu}({P})}\right\rangle,\\ \end{split}$
where we used Lemma A.1 and $d_{\nu}({P})$ is the state visitation distribution starting from $\nu$ and following the policy ${P}$.
The following result, known as Performance Difference Lemma (see e.g. [11, Lemma 1.16]), will be instrumental to prove the convergence rates for PMD in Thm. 7.
@@
Let ${P}_{1}$, ${P}_{2}$ two policy operators. The following equality holds
$J({P}_{1})-J({P}_{2})=\frac{1}{1-\gamma}\left\langle{({P}_{1}-{P}_{2})q({P}_{2})},{d_{\nu}({P}_{1})}\right\rangle.$ (A.9)
@@
Using the definition of $J({P}_{1})$ and Lemma A.1 one gets
$\begin{split}J({P}_{1})-J({P}_{2})&=\left\langle{{P}_{1}\left({Id}-\gamma{T}{P}_{1}\right)^{-1}r},{\nu}\right\rangle-\left\langle{{P}_{2}\left({Id}-\gamma{T}{P}_{2}\right)^{-1}r},{\nu}\right\rangle\\ =&\left\langle{\left({Id}-\gamma{P}_{1}{T}\right)^{-1}{P}_{1}r},{\nu}\right\rangle-\left\langle{{P}_{2}\left({Id}-\gamma{T}{P}_{2}\right)^{-1}r},{\nu}\right\rangle\\ =&\left\langle{\left({Id}-\gamma{P}_{1}{T}\right)^{-1}{P}_{1}\left({Id}-\gamma{T}{P}_{2}\right)\left({Id}-\gamma{T}{P}_{2}\right)^{-1}r},{\nu}\right\rangle\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ -\left\langle{\left({Id}-\gamma{P}_{1}{T}\right)^{-1}\left({Id}-\gamma{P}_{1}{T}\right){P}_{2}\left({Id}-\gamma{T}{P}_{2}\right)^{-1}r},{\nu}\right\rangle\\ =&\left\langle{\left({Id}-\gamma{P}_{1}{T}\right)^{-1}\left[{P}_{1}\left({Id}-\gamma{T}{P}_{2}\right)-\left({Id}-\gamma{P}_{1}{T}\right){P}_{2}\right]\left({Id}-\gamma{T}{P}_{2}\right)^{-1}r},{\nu}\right\rangle\\ =&\left\langle{\left({Id}-\gamma{P}_{1}{T}\right)^{-1}\left[{P}_{1}-{P}_{2}\right]\left({Id}-\gamma{T}{P}_{2}\right)^{-1}r},{\nu}\right\rangle\\ =&\frac{1}{1-\gamma}\left\langle{({P}_{1}-{P}_{2})q({P}_{2})},{d_{\nu}({P}_{1})}\right\rangle.\end{split}$
@@
For any two Markov ${P}_{1},{P}_{2}:B_{b}(\mathcal{X})\to B_{b}(\Omega)$, we have that the directional derivative in ${P}_{1}$ towards ${P}_{2}$ is
$\lim_{h\to 0}\frac{J({P}_{1}+h({P}_{2}-{P}_{1}))-J({P}_{1})}{h}=\frac{1}{1-\gamma}\left\langle{({P}_{2}-{P}_{1})q({P}_{1})},{d_{\nu}({P}_{1})}\right\rangle.$ (A.10)
@@
The result follows by recalling that the space of Markov operators is convex, namely for any $h\in[0,1]$ the term ${P}_{h}={P}_{1}+h({P}_{2}-{P}_{1})$ is still a Markov operator. Therefore, we can apply Lemma A.4 to obtain
$J({P}_{h})-J({P}_{1})$ $=\left\langle{({P}_{h}-{P}_{1})q({P}_{1})},{d_{\nu}({P}_{h})}\right\rangle$ (A.11)
$=h\left\langle{({P}_{2}-{P}_{1})q({P}_{1})},{d_{\nu}({P}_{h})}\right\rangle.$ (A.12)
We can therefore divide the above quantity by $h$ and send $h\to 0$. The result follows by observing that $d_{\nu}({P}_{h})=({Id}-\gamma{T}{P}_{h})^{-*}\nu\to({Id}-\gamma{T}{P}_{1})^{-*}\nu=d_{\nu}({P}_{1})$ for $h\to 0$, since $\left\|{{P}_{h}}\right\|=1$ for any $h\in[0,1]$ and the function ${M}\mapsto(I-\gamma{T}{M})^{-1}$ is continuous on the open ball of radius $11/\gamma>$ 1 in $\mathcal{L}(B_{b}(\Omega),B_{b}(\mathcal{X}))$ with respect to the operator norm. ∎
Properties of $({Id}-\gamma{P}{T})^{-1}$
The quantity $({Id}-\gamma{P}{T})^{-1}$ (note, not $({Id}-\gamma{T}{P})^{-1}$) plays a central role in the study of POWR . We prove here a few properties that will be useful in the following.
Lemma A.6 (Properties of ${Id}-\gamma{P}{T}$).
@@
    For any $f\geq 0\in B_{b}(\mathcal{X})$ it holds $({Id}-\gamma{P}{T})^{-1}f\geq f$.
@@
    The operator $(1-\gamma)({Id}-\gamma{P}{T})^{-1}$ is a Markov operator.
@@
    For any positive measure $\nu\in\mathcal{M}(B_{b}(\mathcal{X}))$ it holds $(1-\gamma)\left\|{({Id}-\gamma{P}{T})^{-*}\nu}\right\|_{{\rm TV}}=\left\|{\nu}\right\|_{{\rm TV}}$.
@@
    For any positive measure $\nu\in\mathcal{M}(B_{b}(\mathcal{X}))$ it holds $\left\|{{P}^{*}\nu}\right\|_{{\rm TV}}=\left\|{\nu}\right\|_{{\rm TV}}$.
@@
    For any bounded linear operator ${X}$, policy operator ${P}$ and discount factor $\gamma<\left\|{{X}}\right\|$, it holds $\bigl{\|}{({Id}-\gamma{X}{P})^{-1}}\bigr{\|}_{\infty}\leq 1/(1-\gamma\left\|{{X}}\right\|)$.
@@
Since both ${T}$ and ${P}$ are Markov operators by construction, it immediately follows that their composition is a Markov operator as well. Using the Neumann series representation of $({Id}-\gamma{P}{T})^{-1}$ it follows that for all $f\geq 0\in B_{b}(\mathcal{X})$
$({Id}-\gamma{P}{T})^{-1}f=\sum_{t=0}^{\infty}\gamma^{t}({P}{T})^{t}f=f+\sum_{t=1}^{\infty}\gamma^{t}({P}{T})^{t}f\geq f\geq 0,$
@@
$({Id}-\gamma{P}{T})^{-1}\mathbf{1}_{\mathcal{X}}=\sum_{t=0}^{\infty}\gamma^{t}({P}{T})^{t}\mathbf{1}_{\mathcal{X}}=\sum_{t=0}^{\infty}\gamma^{t}\mathbf{1}_{\mathcal{X}}=\frac{\mathbf{1}_{\mathcal{X}}}{1-\gamma}$
showing that $(1-\gamma)({Id}-\gamma{P}{T})^{-1}\mathbf{1}_{\mathcal{X}}=\mathbf{1}_{\mathcal{X}}$ and proving (2). Finally, since $(1-\gamma)({Id}-\gamma{P}{T})^{-1}$ and ${P}$ are Markov operators, (3) and (4) follow from the direct application of [25, Theorem 19.2]. For the last point (5), let $f\in B_{b}(\Omega)$. As ${P}$ is a conditional expectation operator, it holds that
$\left\|{{X}{P}f}\right\|_{\infty}\leq\left\|{{X}}\right\|{P}(\mathbf{1}_{\Omega}\left\|{f}\right\|_{\infty})=\left\|{{X}}\right\|\left\|{f}\right\|_{\infty}$
Where the inequality is just the conditional version of Jensen’s inequality [41, Chapter 5] applied on the (convex) $\left\|{\cdot}\right\|_{\infty}$ function, while the equality comes from the fact that ${T}{P}$ is a Markov operator. Then, we have
$\begin{split}\sup_{\left\|{f}\right\|_{\infty}=1}\bigl{\|}{({Id}-\gamma{X}{P})^{-1}f}\bigr{\|}_{\infty}&=\sup_{\left\|{f}\right\|_{\infty}=1}\bigl{\|}{\sum_{t=0}^{\infty}(\gamma{X}{P})^{t}f}\bigr{\|}_{\infty}\\ &\leq\sup_{\left\|{f}\right\|_{\infty}=1}\sum_{t=0}^{\infty}\gamma^{t}\bigl{\|}{({X}{P})^{t}f}\bigr{\|}_{\infty}\\ &\leq\sup_{\left\|{f}\right\|_{\infty}=1}\sum_{t=0}^{\infty}\gamma^{t}\bigl{\|}{{X}}\bigr{\|}^{t}\bigl{\|}{f}\bigr{\|}_{\infty}\\ (\left\|{f}\right\|_{\infty}=1)&=\frac{1}{1-\gamma\bigl{\|}{{X}}\bigr{\|}}.\\ \end{split}$
@@
Let $\gamma>0$ and let ${T}_{1}$, ${T}_{2}$ two linear operators with operator norm strictly less than $\gamma$. Let ${P}$ be a policy operator. Denote by $q({P},{T})=({Id}-\gamma{T}{P})^{-1}r$ the (generalized) action-value function associated to these terms and $v({P},{T})={P}q({P},{T})$ the corresponding value function. Then the following equality holds
$q({P},{T}_{1})-q({P},{T}_{2})=\gamma\left({Id}-\gamma{T}_{1}{P}\right)^{-1}\left({T}_{2}-{T}_{1}\right)v({P},{T}_{2})$ (A.13)
@@
$\begin{split}q({P},{T}_{1})-q({P},{T}_{2})&=\left({Id}-\gamma{T}_{1}{P}\right)^{-1}r-\left({Id}-\gamma{T}_{2}{P}\right)^{-1}r\\ &=\gamma\left({Id}-\gamma{T}_{1}{P}\right)^{-1}\left({T}_{2}-{T}_{1}\right){P}\left({Id}-\gamma{T}_{2}A\right)^{-1}r\\ &=\gamma\left({Id}-\gamma{T}_{1}{P}\right)^{-1}\left({T}_{2}-{T}_{1}\right)v({P},{T}_{2})\end{split}$
where we have used fact that for any two invertible operators ${M}$ and ${P}$ it holds ${M}^{-1}-{P}^{-1}={M}^{-1}({P}-{M}){P}^{-1}$ for the second equation and applied the operatorial characterization of the value function to conclude the proof. ∎
We then have the following result, which hinges a generalization of the standard Simulation lemma in [26, Lemma 2.2] where we account also for the reward function to vary.
@@
Let $\gamma>0$ and let ${T}_{1}$, ${T}_{2}$ two linear operators with operator norm strictly less than $\gamma$. Let $r_{1}$ and $r_{2}$ be two reward functions and ${P}$ a policy operator. Denote by $q({P},{T},r)=({Id}-\gamma{T}{P})^{-1}r$ the (generalized) action-value function associated to these terms and $v({P},{T},r)={P}q({P},{T},r)$ the corresponding value function. Then the following equality holds
$q({P},{T}_{1},r_{1})-q({P},{T}_{2},r_{2})=({Id}-\gamma{T}_{1}{P})^{-1}(r_{1}-r_{2})+\gamma({Id}-\gamma{T}_{1}{P})^{-1}({T}_{2}-{T}_{1})v({P},{T}_{2},r_{2}).$

### v2

@@ A.2 Markov operators and their properties
Recall that since they are Hilbert spaces ${\mathcal{F}}\cong{\mathcal{F}}^{*}$ and ${\mathcal{G}}\cong{\mathcal{G}}^{*}$ are isometric to their dual and therefore we can interpret any $f\in{\mathcal{F}}$ as the function $f(\cdot)=\left\langle{f},{\varphi(\cdot)}\right\rangle$ with some abuse of notation, where clear from context. By Asm. 1 we have that $\mathsf{T}|_{\mathcal{F}}$ takes values in ${\mathcal{G}}$. This means that $(\mathsf{T}|_{\mathcal{F}}f)\in{\mathcal{G}}$ or, in other words
$\left\langle{\mathsf{T}|_{\mathcal{F}}f},{\psi(x,a)}\right\rangle$ $=(\mathsf{T}|_{\mathcal{F}}f)(x,a)$
@@
$\left\langle{f},{(\mathsf{T}|_{\mathcal{F}})^{*}\psi(x,a)}\right\rangle=\left\langle{f},{\int\varphi(x^{\prime})\leavevmode\nobreak\ \tau(dx^{\prime}|x,a)}\right\rangle.$
@@
We note that the result can be extended to the setting where $\mathsf{T}|_{{\mathcal{F}}}({\mathcal{F}})\subseteq{\mathcal{G}}$, namely the image of $\mathsf{T}|_{{\mathcal{F}}}$ is contained in ${\mathcal{G}}$, namely a sort of $({\mathcal{F}},{\mathcal{G}})$-compatibility for the transition operator (see Def. 1).
@@ A.3 The operatorial formulation of RL
According to the operatorial characterization in 7, the action value function of a policy $\pi$ is directly related to the action of the associated policy operator $\mathsf{P}$. To highlight this relation, we will adopt the following notation:
@@
    $q(\mathsf{P})=\left(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}\right)^{-1}r.$ (A.5)
@@
    $J(\mathsf{P})=\left\langle{\mathsf{P}\left(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}\right)^{-1}r},{\nu}\right\rangle=\left\langle{\mathsf{P}q(\mathsf{P})},{\nu}\right\rangle=\left\langle{v(\mathsf{P})},{\nu}\right\rangle.$ (A.7)
  • State visitation (or State occupancy) measure. By the characterization of the adjoints of $\mathsf{P}$ and $\mathsf{T}$ (see discussion in Sec. 3 we can represent the evolution of a state distribution $\nu_{t}$ at time $t$ to the next state distribution as $\nu_{t+1}=\mathsf{T}^{*}\mathsf{P}^{*}\nu_{t}$. Applying this relation recursively, we recover the state visitation probability associated to the starting state distribution $\nu_{0}=\nu\in\mathcal{P}(\mathcal{X})$, the MDP with transition $\mathsf{T}$ and the policy $\mathsf{P}$ as
    $d_{\nu}(\mathsf{P})=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}(\mathsf{T}^{*}\mathsf{P}^{*})^{t}\nu=(1-\gamma)\left(\mathsf{Id}-\gamma\mathsf{P}\mathsf{T}\right)^{-*}\nu,$ (A.8)
    where the $(1-\gamma)\gamma^{t}$ is a normalizing factor to guarantee that the series corresponds to a convex combination of the probability distributions $\nu_{t}$, hence guaranteeing $d_{\nu}(\mathsf{P})$ to be well-defined (namely it belongs to $\mathcal{P}(\mathcal{X})$).
@@
Algebraic manipulation of the cumulative expected reward $J(\mathsf{P})$ implies
$\begin{split}J(\mathsf{P})&=\left\langle{\mathsf{P}\left(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}\right)^{-1}r},{\nu}\right\rangle=\left\langle{\mathsf{P}r},{\left(\mathsf{Id}-\gamma\mathsf{P}\mathsf{T}\right)^{-*}\nu}\right\rangle=\frac{1}{1-\gamma}\left\langle{\mathsf{P}r},{d_{\nu}(\mathsf{P})}\right\rangle,\\ \end{split}$
where we used Lemma A.1 and $d_{\nu}(\mathsf{P})$ is the state visitation distribution starting from $\nu$ and following the policy $\mathsf{P}$.
The following result, known as Performance Difference Lemma [[, see e.g.]Lemma 1.16]agarwal2021, will be instrumental to prove the convergence rates for PMD in Thm. 7.
@@
Let $\mathsf{P}_{1}$, $\mathsf{P}_{2}$ two policy operators. The following equality holds
$J(\mathsf{P}_{1})-J(\mathsf{P}_{2})=\frac{1}{1-\gamma}\left\langle{(\mathsf{P}_{1}-\mathsf{P}_{2})q(\mathsf{P}_{2})},{d_{\nu}(\mathsf{P}_{1})}\right\rangle.$ (A.9)
@@
Using the definition of $J(\mathsf{P}_{1})$ and Lemma A.1 one gets
$\begin{split}J(\mathsf{P}_{1})-J(\mathsf{P}_{2})&=\left\langle{\mathsf{P}_{1}\left(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{1}\right)^{-1}r},{\nu}\right\rangle-\left\langle{\mathsf{P}_{2}\left(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{2}\right)^{-1}r},{\nu}\right\rangle\\ =&\left\langle{\left(\mathsf{Id}-\gamma\mathsf{P}_{1}\mathsf{T}\right)^{-1}\mathsf{P}_{1}r},{\nu}\right\rangle-\left\langle{\mathsf{P}_{2}\left(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{2}\right)^{-1}r},{\nu}\right\rangle\\ =&\left\langle{\left(\mathsf{Id}-\gamma\mathsf{P}_{1}\mathsf{T}\right)^{-1}\mathsf{P}_{1}\left(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{2}\right)\left(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{2}\right)^{-1}r},{\nu}\right\rangle\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ -\left\langle{\left(\mathsf{Id}-\gamma\mathsf{P}_{1}\mathsf{T}\right)^{-1}\left(\mathsf{Id}-\gamma\mathsf{P}_{1}\mathsf{T}\right)\mathsf{P}_{2}\left(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{2}\right)^{-1}r},{\nu}\right\rangle\\ =&\left\langle{\left(\mathsf{Id}-\gamma\mathsf{P}_{1}\mathsf{T}\right)^{-1}\left[\mathsf{P}_{1}\left(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{2}\right)-\left(\mathsf{Id}-\gamma\mathsf{P}_{1}\mathsf{T}\right)\mathsf{P}_{2}\right]\left(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{2}\right)^{-1}r},{\nu}\right\rangle\\ =&\left\langle{\left(\mathsf{Id}-\gamma\mathsf{P}_{1}\mathsf{T}\right)^{-1}\left[\mathsf{P}_{1}-\mathsf{P}_{2}\right]\left(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{2}\right)^{-1}r},{\nu}\right\rangle\\ =&\frac{1}{1-\gamma}\left\langle{(\mathsf{P}_{1}-\mathsf{P}_{2})q(\mathsf{P}_{2})},{d_{\nu}(\mathsf{P}_{1})}\right\rangle.\end{split}$
@@
For any two Markov $\mathsf{P}_{1},\mathsf{P}_{2}:B_{b}(\mathcal{X})\to B_{b}(\Omega)$, we have that the directional derivative in $\mathsf{P}_{1}$ towards $\mathsf{P}_{2}$ is
$\lim_{h\to 0}\frac{J(\mathsf{P}_{1}+h(\mathsf{P}_{2}-\mathsf{P}_{1}))-J(\mathsf{P}_{1})}{h}=\frac{1}{1-\gamma}\left\langle{(\mathsf{P}_{2}-\mathsf{P}_{1})q(\mathsf{P}_{1})},{d_{\nu}(\mathsf{P}_{1})}\right\rangle.$ (A.10)
@@
The result follows by recalling that the space of Markov operators is convex, namely for any $h\in[0,1]$ the term $\mathsf{P}_{h}=\mathsf{P}_{1}+h(\mathsf{P}_{2}-\mathsf{P}_{1})$ is still a Markov operator. Therefore, we can apply Lemma A.3 to obtain
$J(\mathsf{P}_{h})-J(\mathsf{P}_{1})$ $=\left\langle{(\mathsf{P}_{h}-\mathsf{P}_{1})q(\mathsf{P}_{1})},{d_{\nu}(\mathsf{P}_{h})}\right\rangle$ (A.11)
$=h\left\langle{(\mathsf{P}_{2}-\mathsf{P}_{1})q(\mathsf{P}_{1})},{d_{\nu}(\mathsf{P}_{h})}\right\rangle.$ (A.12)
We can therefore divide the above quantity by $h$ and send $h\to 0$. The result follows by observing that $d_{\nu}(\mathsf{P}_{h})=(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{h})^{-*}\nu\to(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{1})^{-*}\nu=d_{\nu}(\mathsf{P}_{1})$ for $h\to 0$, since $\left\|{\mathsf{P}_{h}}\right\|=1$ for any $h\in[0,1]$ and the function $\mathsf{M}\mapsto(I-\gamma\mathsf{T}\mathsf{M})^{-1}$ is continuous on the open ball of radius $11/\gamma>$ 1 in $\mathcal{L}(B_{b}(\Omega),B_{b}(\mathcal{X}))$ with respect to the operator norm. ∎
Properties of $(\mathsf{Id}-\gamma\mathsf{P}\mathsf{T})^{-1}$
The quantity $(\mathsf{Id}-\gamma\mathsf{P}\mathsf{T})^{-1}$ (note, not $(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P})^{-1}$) plays a central role in the study of POWR. We prove here a few properties that will be useful in the following.
Lemma A.5 (Properties of $\mathsf{Id}-\gamma\mathsf{P}\mathsf{T}$).
@@
    For any $f\geq 0\in B_{b}(\mathcal{X})$ it holds $(\mathsf{Id}-\gamma\mathsf{P}\mathsf{T})^{-1}f\geq f$.
@@
    The operator $(1-\gamma)(\mathsf{Id}-\gamma\mathsf{P}\mathsf{T})^{-1}$ is a Markov operator.
@@
    For any positive measure $\nu\in\mathcal{M}(B_{b}(\mathcal{X}))$ it holds $(1-\gamma)\left\|{(\mathsf{Id}-\gamma\mathsf{P}\mathsf{T})^{-*}\nu}\right\|_{{\rm TV}}=\left\|{\nu}\right\|_{{\rm TV}}$.
@@
    For any positive measure $\nu\in\mathcal{M}(B_{b}(\mathcal{X}))$ it holds $\left\|{\mathsf{P}^{*}\nu}\right\|_{{\rm TV}}=\left\|{\nu}\right\|_{{\rm TV}}$.
@@
    For any bounded linear operator $\mathsf{X}$, policy operator $\mathsf{P}$ and discount factor $\gamma<\left\|{\mathsf{X}}\right\|$, it holds $\bigl{\|}{(\mathsf{Id}-\gamma\mathsf{X}\mathsf{P})^{-1}}\bigr{\|}_{\infty}\leq 1/(1-\gamma\left\|{\mathsf{X}}\right\|)$.
@@
Since both $\mathsf{T}$ and $\mathsf{P}$ are Markov operators by construction, it immediately follows that their composition is a Markov operator as well. Using the Neumann series representation of $(\mathsf{Id}-\gamma\mathsf{P}\mathsf{T})^{-1}$ it follows that for all $f\geq 0\in B_{b}(\mathcal{X})$
$(\mathsf{Id}-\gamma\mathsf{P}\mathsf{T})^{-1}f=\sum_{t=0}^{\infty}\gamma^{t}(\mathsf{P}\mathsf{T})^{t}f=f+\sum_{t=1}^{\infty}\gamma^{t}(\mathsf{P}\mathsf{T})^{t}f\geq f\geq 0,$
@@
$(\mathsf{Id}-\gamma\mathsf{P}\mathsf{T})^{-1}\mathbf{1}_{\mathcal{X}}=\sum_{t=0}^{\infty}\gamma^{t}(\mathsf{P}\mathsf{T})^{t}\mathbf{1}_{\mathcal{X}}=\sum_{t=0}^{\infty}\gamma^{t}\mathbf{1}_{\mathcal{X}}=\frac{\mathbf{1}_{\mathcal{X}}}{1-\gamma}$
showing that $(1-\gamma)(\mathsf{Id}-\gamma\mathsf{P}\mathsf{T})^{-1}\mathbf{1}_{\mathcal{X}}=\mathbf{1}_{\mathcal{X}}$ and proving (2). Finally, since $(1-\gamma)(\mathsf{Id}-\gamma\mathsf{P}\mathsf{T})^{-1}$ and $\mathsf{P}$ are Markov operators, (3) and (4) follow from the direct application of [29, Theorem 19.2]. For the last point (5), let $f\in B_{b}(\Omega)$. As $\mathsf{P}$ is a conditional expectation operator, it holds that
$\left\|{\mathsf{X}\mathsf{P}f}\right\|_{\infty}\leq\left\|{\mathsf{X}}\right\|\mathsf{P}(\mathbf{1}_{\Omega}\left\|{f}\right\|_{\infty})=\left\|{\mathsf{X}}\right\|\left\|{f}\right\|_{\infty}$
Where the inequality is just the conditional version of Jensen’s inequality [44, Chapter 5] applied on the (convex) $\left\|{\cdot}\right\|_{\infty}$ function, while the equality comes from the fact that $\mathsf{T}\mathsf{P}$ is a Markov operator. Then, we have
$\begin{split}\sup_{\left\|{f}\right\|_{\infty}=1}\bigl{\|}{(\mathsf{Id}-\gamma\mathsf{X}\mathsf{P})^{-1}f}\bigr{\|}_{\infty}&=\sup_{\left\|{f}\right\|_{\infty}=1}\bigl{\|}{\sum_{t=0}^{\infty}(\gamma\mathsf{X}\mathsf{P})^{t}f}\bigr{\|}_{\infty}\\ &\leq\sup_{\left\|{f}\right\|_{\infty}=1}\sum_{t=0}^{\infty}\gamma^{t}\bigl{\|}{(\mathsf{X}\mathsf{P})^{t}f}\bigr{\|}_{\infty}\\ &\leq\sup_{\left\|{f}\right\|_{\infty}=1}\sum_{t=0}^{\infty}\gamma^{t}\bigl{\|}{\mathsf{X}}\bigr{\|}^{t}\bigl{\|}{f}\bigr{\|}_{\infty}\\ (\left\|{f}\right\|_{\infty}=1)&=\frac{1}{1-\gamma\bigl{\|}{\mathsf{X}}\bigr{\|}}.\\ \end{split}$
@@
Let $\gamma>0$ and let $\mathsf{T}_{1}$, $\mathsf{T}_{2}$ two linear operators with operator norm strictly less than $\gamma$. Let $\mathsf{P}$ be a policy operator. Denote by $q(\mathsf{P},\mathsf{T})=(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P})^{-1}r$ the (generalized) action-value function associated to these terms and $v(\mathsf{P},\mathsf{T})=\mathsf{P}q(\mathsf{P},\mathsf{T})$ the corresponding value function. Then the following equality holds
$q(\mathsf{P},\mathsf{T}_{1})-q(\mathsf{P},\mathsf{T}_{2})=\gamma\left(\mathsf{Id}-\gamma\mathsf{T}_{1}\mathsf{P}\right)^{-1}\left(\mathsf{T}_{2}-\mathsf{T}_{1}\right)v(\mathsf{P},\mathsf{T}_{2})$ (A.13)
@@
$\begin{split}q(\mathsf{P},\mathsf{T}_{1})-q(\mathsf{P},\mathsf{T}_{2})&=\left(\mathsf{Id}-\gamma\mathsf{T}_{1}\mathsf{P}\right)^{-1}r-\left(\mathsf{Id}-\gamma\mathsf{T}_{2}\mathsf{P}\right)^{-1}r\\ &=\gamma\left(\mathsf{Id}-\gamma\mathsf{T}_{1}\mathsf{P}\right)^{-1}\left(\mathsf{T}_{2}-\mathsf{T}_{1}\right)\mathsf{P}\left(\mathsf{Id}-\gamma\mathsf{T}_{2}A\right)^{-1}r\\ &=\gamma\left(\mathsf{Id}-\gamma\mathsf{T}_{1}\mathsf{P}\right)^{-1}\left(\mathsf{T}_{2}-\mathsf{T}_{1}\right)v(\mathsf{P},\mathsf{T}_{2})\end{split}$
where we have used fact that for any two invertible operators $\mathsf{M}$ and $\mathsf{P}$ it holds $\mathsf{M}^{-1}-\mathsf{P}^{-1}=\mathsf{M}^{-1}(\mathsf{P}-\mathsf{M})\mathsf{P}^{-1}$ for the second equation and applied the operatorial characterization of the value function to conclude the proof. ∎
We then have the following result, which hinges on a generalization of the standard Simulation lemma in [30, Lemma 2.2] where we account also for the reward function to vary.
@@
Let $\gamma>0$ and let $\mathsf{T}_{1}$, $\mathsf{T}_{2}$ two linear operators with operator norm strictly less than $\gamma$. Let $r_{1}$ and $r_{2}$ be two reward functions and $\mathsf{P}$ a policy operator. Denote by $q(\mathsf{P},\mathsf{T},r)=(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P})^{-1}r$ the (generalized) action-value function associated to these terms and $v(\mathsf{P},\mathsf{T},r)=\mathsf{P}q(\mathsf{P},\mathsf{T},r)$ the corresponding value function. Then the following equality holds
$q(\mathsf{P},\mathsf{T}_{1},r_{1})-q(\mathsf{P},\mathsf{T}_{2},r_{2})=(\mathsf{Id}-\gamma\mathsf{T}_{1}\mathsf{P})^{-1}(r_{1}-r_{2})+\gamma(\mathsf{Id}-\gamma\mathsf{T}_{1}\mathsf{P})^{-1}(\mathsf{T}_{2}-\mathsf{T}_{1})v(\mathsf{P},\mathsf{T}_{2},r_{2}).$